# Enhancing Multimodal Medical Image Classification using Cross-Graph Modal Contrastive Learning

### v4

@@ Abstract
The classification of medical images is a pivotal aspect of disease diagnosis, often enhanced by deep learning techniques. However, traditional approaches typically focus on unimodal medical image data, neglecting the integration of diverse non-image patient data. This paper proposes a novel Cross-Graph Modal Contrastive Learning (CGMCL) framework for multimodal structured data from different data domains to improve medical image classification. The model effectively integrates both image and non-image data by constructing cross-modality graphs and leveraging contrastive learning to align multimodal features in a shared latent space. An inter-modality feature scaling module further optimizes the representation learning process by reducing the gap between heterogeneous modalities. The proposed approach is evaluated on two datasets: a Parkinson’s disease (PD) dataset and a public melanoma dataset. Results demonstrate that CGMCL outperforms conventional unimodal methods in accuracy, interpretability, and early disease prediction. Additionally, the method shows superior performance in multi-class melanoma classification. The CGMCL framework provides valuable insights into medical image classification while offering improved disease interpretability and predictive capabilities. The code and data are available from https://github.com/Ding1119/CGMCL
Index Terms:
@@
Neurodegenerative disease, SPECT, Contrastive learning, Multimodal fusion, Classification, Cross-graph modal graph learning.
I Introduction
@@
In recent years, medical computing has increasingly shifted toward utilizing deep learning frameworks as a foundation for diagnostic analysis. In particular, multimodal deep learning has emerged as a powerful approach for integrating diverse data sources to enhance predictive accuracy in various medical applications. The complex and heterogeneous nature of medical data, including radiological images, clinical records and histological images, necessitates the use of multimodal fusion methods to extract complementary features and improve clinical decision-making [1, 2, 3]. However, in real-world clinical settings, many modalities show inconsistency in modality distributions depending on disease characteristics, which indirectly creates limitations for multimodal learning [4]. Addressing these challenges, some studies have proposed advanced fusion architectures, such as feature-level fusion, decision-level fusion, and hybrid fusion, to optimize the integration of multimodal information [2, 5]. For example, integrating dermatological images with patient metadata significantly enhances the classification accuracy of skin cancer detection systems, outperforming unimodal image-based models [5].
@@
Graph-based deep learning such as graph convolutional neural networks (GCNs) [6] learn from unstructured features by processing non-Euclidean distances, enabling them to perform various downstream tasks, including node classification, link prediction, and graph classification. An advanced framework of GCNs offer a promising avenue for identifying commonalities among patients in disease relationships via medical imaging data or symptom similarity [7, 8]. However, these methods only perform disease prediction on a single modality and lack information from both images and meta-features. AMA-GCN [9] uses multi-layer aggregation mechanisms and automatic feature selection to integrate both image and non-image data for various diseases. Nevertheless, AMA-GCN’s thresholds and aggregation methods require readjustment, and the model may overly depend on certain features, potentially leading to overfitting. Moreover, there is no unified approach for effectively achieving clinical interpretability through multimodal fusion, particularly for images and metadata associated with distinct graph structures in clinical settings, which should be further leveraged to enhance efficient latent space learning [10].
@@
Recent studies have focused on aligning multimodal data in a shared embedding space for Alzheimer’s disease (AD) prediction. Notably, attention mechanisms applied to tabular data have demonstrated improved model performance and interpretability [11], while multimodal transformers combining image and clinical data have shown success in predicting AD progression [12]. Furthermore,  [13] has demonstrated the effectiveness of self-supervised contrastive learning in handling different modalities through dedicated encoder networks, which map data into a shared feature space while optimizing a contrastive loss function. However, these studies have primarily focused on specific diseases, with limited exploration of model generalization across diverse multimodal disease datasets and a lack of meta-features interpretable for disease classification tasks.
@@
To address above problems, we proposes a novel approach that leverages dual graph networks to construct a cross-graph modal feature topology for diverse modalities, such as medical images and quantitative parameters. Our method employs graph attention networks (GATs) [14] to learn feature encodings across modalities through node embedding. Additionally, we design a graph contrastive loss to enhance the similarity of the final representations. We validate our proposed Cross-Graph Modal Contrastive Learning (CGMCL) approach for multimodal fusion, and clinical interpretability using two multimodal medical datasets: a private PD dataset and a public melanoma dataset. The main contributions of this study can be summarized as follows:
  • We propose a multimodal fusion framework incorporating multi-structured modalities (e.g., medical images and metadata) from different patient disease characteristics as graph-structured features to enhance classification accuracy through scale and eliminate multimodality distribution inconsistent gaps.
@@
  • To better evaluate the robustness of our proposed CGMCL model, we achieved more effective performance compared to fashion CNN-based models across two different types of multimodal datasets.
@@
II RELATED WORK
II-A Multimodal Learning in in Medical Imaging
@@
Traditional image-based models approaches often fail to capture the full complexity of diseases, whereas multimodal fusion leverages complementary information to enhance predictive performance [15]. Some non-deep learning research works focus on using Wavelet Transform for multimodal image fusion in MRI/CT, PET/CT, and PET/MRI [16, 17, 18]. In neurodegenerative diseases diagnosis, studies have demonstrated that combining T1-weighted MRI, PET, and cerebrospinal fluid (CSF) biomarkers significantly improves early detection and progression monitoring of Parkinson’s Disease (PD) and Alzheimer’s Disease (AD) [19, 15]. Moreover, recent multimodal deep learning research focuses on PD classification applied multi-co-attention model successfully integrated brain single-photon emission computed tomography (SPECT) images with DNA methylation data [20]. However, in the above studies, feature extraction from medical images has predominantly been conducted at the pixel level, with limited consideration for structured feature learning that accounts for inter-patient image similarities (e.g., graph structures). Therefore, we propose that early diagnosis, particularly for conditions such as PD, requires more precise multimodal models to improve early prediction and clinical interpretability.
II-B Multi-Modality Fusion Strategies
According to deep learning fusion strategies, including feature extraction from two modalities, followed by fusion approaches such as vector concatenation, attention-based fusion, and joint network fusion learning to combine cross-domain features [21, 22, 23, 24, 21], as shown in Fig. 1 (a)-(c). Recent advancements in complicated multimodal framework have introduced more sophisticated fusion techniques, such as transformer-based fusion. MATR has demonstrated superior performance in multimodal medical image fusion by leveraging adaptive convolution and transformer mechanisms to utilize local and global contextual information [25]. However, the greatest challenge with the approach mentioned above is that it does not consider the different characteristics of different modalities when performing modality fusion, which creates gaps in feature representation between their distributions.
To learn consistent latent representations, we designed a dual-graph architecture. The first graph utilizes patient image features extracted from a pre-trained ResNet model as node representations, where each node represents a patient. The second graph incorporates structured meta-features, as illustrated in Fig. 1 (d). We then employ a dual-structured graph attention-based fusion framework to project and scale these different modalities into a consistent feature space, as shown in Fig. 2
@@ II-B Multi-Modality Fusion Strategies
III METHODOLOGY
III-A Problem Definition and Formulation
In this section, we formulate similarities within a single modality across different patients as graph structures to enhance disease classification compared to unimodal approaches. This study focuses on two medical modalities: medical images and non-image data (e.g., meta-features). We denote a multimodal set $\textbf{X}=\{X_{1},X_{2},...,X_{N}\}$ for $N$ patients. We can represent the data for the $i$-th patient as a 3-tuple $X_{i}=\{(I_{i},C_{i},Y_{i})\}_{i=1}^{N}$, where $I_{i}\in\mathbb{R}^{h\times h}$ represents the medical image with a scale dimension $h$, $C_{i}\in\mathbb{R}^{F}$ represents the meta-features with $F$ features, and $Y_{i}$ corresponds to the disease labels. We first construct a non-linear model (e.g., an CNN) to generate initial feature maps $I_{i}^{\prime}$ from the images:
@@ III-A Problem Definition and Formulation
where the function $f(\cdot):\mathbb{R}^{D}\to\mathbb{R}^{d}$ represents the feature extractor for images, $D$ is the dimension of the initial latent space in the $l-1$ layer, and $d$ is the reduced feature dimension after the $l$-th convolution layer.
III-B Graph Construction
In practice, to better measure the structural properties of the two modalities for each patient in non-Euclidean distance feature spaces, we consider the two modality graphs, structured-image graph $\mathcal{G}^{I}(\mathcal{E}^{I},\mathcal{V}^{I})$ and structured meta-feature graph $\mathcal{G}^{C}(\mathcal{E}^{C},\mathcal{V}^{C})$, with edges $\left|\mathcal{E}^{I}\right|$ and $\left|\mathcal{E}^{C}\right|$ and vertices $\left|\mathcal{V}^{I}\right|$ and $\left|\mathcal{V}^{C}\right|$, respectively. Given the input encoded features $F_{i}$ (e.g., $I_{i}^{\prime}$ or $C_{i}$), we can construct the binary adjacency matrices $A^{I}$ and $A^{C}$ using a $K$-nearest neighbors graph [26]:
[…]
@@ III-B Graph Construction
III-C Graph Attention Encoder
Our modality-based graph model employs a GAT as the foundational features encoder to facilitate cross-graph modal learning across different graph modalities. Given the two input features node representations $I^{\prime}=[I^{\prime}_{1},I^{\prime}_{2},...,I^{\prime}_{N}]\in\mathbb{R}^{N\times d}$ and $C=[C_{1},C_{2},...,C_{N}]\in\mathbb{R}^{N\times F}$, we can formulate the multi-modality attention coefficients in a GNN as:
@@ III-C Graph Attention Encoder
$\vspace{-1mm}H_{i}^{I^{\prime}}=\sigma\left(\sum_{j\in\mathcal{N}_{i}}\alpha^{I^{\prime}}_{ij}\mathbf{W}^{I^{\prime}}{I^{\prime}}_{j}\right),$ (7)
@@
We focus on enhancing the fusion of learned representations across cross-graph node modality embeddings. To integrate both the GAT output representations $H_{i}^{I^{\prime}}$ and $H_{i}^{C}$ with the encoder feature information, we also incorporate a concatenation of features generated by the CNN encoder $I^{\prime}$ and the meta-features $C$ into the graph encoder. This can be expressed as:
$\mathcal{H}^{I}=\left[H_{i}^{I}\mathbin{\|}I_{i}^{\prime}\right],$ (9)
@@
[…]
III-D Inter-Modality Feature Enhancement and Scaling
To address scale discrepancies between features and reduce modality gaps in the latent space for two single-modality output features, we introduce the inter-modality feature enhancement and scaling (IMFES) module. This module is designed to enhance the learning of intrinsic modality distributions and preserve important structural information within each modality. The process begins by applying a simple multilayer perceptron (MLP) to the concatenated tensors $\mathcal{H}^{I}$ and $\mathcal{H}^{C}$, transforming them into a single-modality probability matrix. Next, we perform element-wise multiplication between the GAT encoder outputs $H^{I^{\prime}}$ and $H^{C}$, and the non-linear MLP transforms $\zeta(\mathcal{H}^{I})$ and $\zeta(\mathcal{C}^{I})$, where $\zeta(\cdot)=\mathrm{MLP(\cdot)}$. This operation enables fine-grained extraction of individual modality features, as represented by the following equations:

### v5

# Enhancing Multimodal Medical Image Classification through Cross-Graph Modal Contrastive Learning

@@ Abstract
The classification of medical images is a pivotal aspect of disease diagnosis, often enhanced by deep learning techniques. However, traditional approaches typically focus on unimodal medical image data, neglecting the integration of diverse non-image patient data. This paper proposes a novel Cross-Graph Modal Contrastive Learning (CGMCL) framework for multimodal structured data from different data domains to improve medical image classification. The model effectively integrates both image and non-image data by constructing cross-modality graphs and leveraging contrastive learning to align multimodal features in a shared latent space. An inter-modality feature scaling module further optimizes the representation learning process by reducing the gap between heterogeneous modalities. The proposed approach is evaluated on two datasets: a Parkinson’s disease (PD) dataset and a public melanoma dataset. Results demonstrate that CGMCL outperforms conventional unimodal methods in accuracy, interpretability, and early disease prediction. Additionally, the method shows superior performance in multi-class melanoma classification. The CGMCL framework provides valuable insights into medical image classification while offering improved disease interpretability and predictive capabilities. The code and data are available from https://github.com/Ding1119/CGMCL.
keywords:
@@
Neurodegenerative disease, Single Photon Emission Computed Tomography, Contrastive learning, Multimodal fusion, Classification, Parkinson’s disease, Graph Neural Networks
journal: Expert Systems with Applications
1 Introduction
@@
In recent years, medical computing has increasingly shifted toward utilizing deep learning frameworks as a foundation for diagnostic analysis. In particular, multimodal deep learning has emerged as a powerful approach for integrating diverse data sources to enhance predictive accuracy in various medical applications. The complex and heterogeneous nature of medical data, including radiological images, clinical records and histological images, necessitates the use of multimodal fusion methods to extract complementary features and improve clinical decision-making yang2022large, nakach2024comprehensive, rupp2023exbehrt. However, in real-world clinical settings, many modalities show inconsistency in modality distributions depending on disease characteristics, which indirectly creates limitations for multimodal learning zhang2024multimodal. Addressing these challenges, some studies have proposed advanced fusion architectures, such as feature-level fusion, decision-level fusion, and hybrid fusion, to optimize the integration of multimodal information nakach2024comprehensive, tang2024joint. For example, integrating dermatological images with patient metadata significantly enhances the classification accuracy of skin cancer detection systems, outperforming unimodal image-based models tang2024joint.
@@
Graph-based deep learning such as graph convolutional neural networks (GCNs) kipf2017semi learn from unstructured features by processing non-Euclidean distances, enabling them to perform various downstream tasks, including node classification, link prediction, and graph classification. An advanced framework of GCNs offer a promising avenue for identifying commonalities among patients in disease relationships via medical imaging data or symptom similarity lu2021weighted, mao2022imagegcn. However, these methods only perform disease prediction on a single modality and lack information from both images and meta-features. AMA-GCN chen2021ama uses multi-layer aggregation mechanisms and automatic feature selection to integrate both image and non-image data for various diseases. Nevertheless, AMA-GCN’s thresholds and aggregation methods require readjustment, and the model may overly depend on certain features, potentially leading to overfitting. Moreover, there is no unified approach for effectively achieving clinical interpretability through multimodal fusion, particularly for images and metadata associated with distinct graph structures in clinical settings, which should be further leveraged to enhance efficient latent space learning wang2020gcn.
@@
Recent studies have focused on aligning multimodal data in a shared embedding space for Alzheimer’s disease (AD) prediction. Notably, attention mechanisms applied to tabular data have demonstrated improved model performance and interpretability huang2023multimodal, while multimodal transformers combining image and clinical data have shown success in predicting AD progression liu2023triformer. Furthermore,  taleb2022contig has demonstrated the effectiveness of self-supervised contrastive learning in handling different modalities through dedicated encoder networks, which map data into a shared feature space while optimizing a contrastive loss function.
However, existing studies have primarily focused on specific diseases, with limited investigation into model generalization across diverse multimodal disease datasets and a lack of interpretable meta-features for disease classification. Recent approaches remain largely tailored to single-disease contexts and struggle to generalize across heterogeneous multimodal medical datasets that differ in data structure, modality distribution, and acquisition sources teoh2024advancing, li2024review. This heterogeneity, which arises from intrinsic disparities among medical images, quantitative biomarkers, and clinical metadata, presents substantial challenges for achieving robust multimodal fusion and interpretability liu2025review.
@@
To address these limitations, we propose a novel Cross-Graph Modal Contrastive Learning (CGMCL) framework that explicitly models multimodal heterogeneity through a dual-graph architecture. CGMCL constructs a cross-graph modal feature topology that represents diverse modalities, such as imaging data and quantitative parameters, as structured graphs. Using graph attention networks (GATs) Velickovic_2017, our method learns node-level embeddings that capture inter-modal relationships, while a graph contrastive loss enforces cross-graph consistency and improves heterogeneous modality representation alignment. We validate the proposed CGMCL framework on two heterogeneous multimodal medical datasets, including a private PD dataset and a public melanoma dataset, to demonstrate its fusion capability and clinical interpretability. The main contributions of this study can be summarized as follows:
  • 1.
    We introduce a multimodal fusion framework that integrates multi-structured and heterogeneous modalities (e.g., medical images and meta-features) by representing patient characteristics as graph-structured features, thereby mitigating modality distribution inconsistencies and improving classification performance.
  • 2.
@@
    To better evaluate the robustness of our proposed CGMCL model, we achieved more effective performance compared to fashion CNN-based models across two distinct types of multimodal datasets.
  • 3.
@@
  • 4.
    We establish the first large-scale multimodal annotation dataset for early-stage PD patients and leverage CGMCL to quantify the relative importance of early degenerative symptoms, enabling better identification of early PD markers.
2 RELATED WORK
2.1 SPECT for Parkinsonism Diagnosis
Dopamine imbalance generates toxic metabolites and oxidative stress, driving dopaminergic neuron degeneration and Parkinson’s disease bucher2020acquired, zhou2023role. To assess the integrity of dopaminergic neurons, Dopamine transporter (DAT) imaging with single photon emission computed tomography (SPECT) can detect presynaptic dopaminergic neuronal dysfunction, serving as a valuable tool for differentiating PD from conditions without presynaptic dopaminergic deficit brogley2019datquant. For interpretation of DAT-SPECT, automatic semi-quantification of DAT-SPECT images with DaTQUANT package (GE Healthcare) provides objective parameters for assistance of image interpretation as shown in Fig. 1.
Existing research has developed computer-aided diagnosis (CAD) systems for PD using DaTSCAN SPECT, which employ machine learning to analyze striatal dopamine transporter deficits. The approaches include high-dimensional voxel-based methods with dimensionality reduction segovia2012improved, rojas2013application and simplified striatal binding ratio (SBR) features that achieve accuracies exceeding 9696prashanth2014automatic. Among these, putamen SBR demonstrates superior discriminative power compared to caudate SBR, while logistic regression further enables risk probability estimation for patient stratification. However, single-modality SBR-based methods may not reflect the heterogeneous presentations encountered in routine clinical practice serag2025multimodal, zhu2024multimodal, particularly as high-dimensional voxel-based approaches require extensive feature reduction that potentially discards discriminative information solana2021classification.
Figure 1: The semi-quantification of parameters was derived from the DaTQUANT package and described the location of the striatum, relating to the entire striatum, caudate nucleus, entire putamen, anterior putamen, and posterior putamen takatsu2023dysfunction.
2.2 Multimodal Learning in Medical Imaging
@@
Traditional image-based models approaches often fail to capture the full complexity of diseases, whereas multimodal fusion leverages complementary information to enhance predictive performance pahuja2022deep. Some non-deep learning research works focus on using Wavelet Transform for multimodal image fusion in MRI/CT, PET/CT, and PET/MRI zheng2007new, cheng2008medical, bhavana2015multi. In neurodegenerative diseases diagnosis, studies have demonstrated that combining T1-weighted MRI, PET, and cerebrospinal fluid (CSF) biomarkers significantly improves early detection and progression monitoring of Parkinson’s Disease (PD) and Alzheimer’s Disease (AD) dentamaro2024enhancing, pahuja2022deep.
Recent research has introduced multimodal approaches for PD diagnosis and progression assessment. While traditional deep learning studies based on single modalities such as speech or imaging demonstrate useful discriminative power, they face limitations in generalizability ul2022survey. To address this limitation,  lei2017joint proposed a multimodal sparse learning framework that integrates magnetic resonance imaging (MRI) and diffusion tensor imaging (DTI) to simultaneously perform disease classification and clinical score prediction, demonstrating the robustness of multimodal fusion in PD diagnosis. Moreover, recent multimodal deep learning research focuses on PD classification applied multi-co-attention model successfully integrated brain single-photon emission computed tomography (SPECT) images with DNA methylation data Taylor_2019. However, in the above studies, feature extraction from medical images has predominantly been conducted at the pixel level, with limited consideration for structured feature learning that accounts for inter-patient image similarities (e.g., graph structures). Therefore, we propose that early diagnosis, particularly for conditions such as PD, requires more precise multimodal models to improve early prediction and clinical interpretability.
@@
2.3 Multi-Modality Fusion Strategies
According to deep learning fusion strategies, including feature extraction from two modalities, followed by fusion approaches such as vector concatenation, attention-based fusion, and joint network fusion learning to combine cross-domain features cui2023deep, li2024review, duan2024deep, stahlschmidt2022multimodal, cui2023deep, as shown in Fig. 2 (a)-(c). Recent advancements in complicated multimodal framework have introduced more sophisticated fusion techniques, such as transformer-based fusion. MATR has demonstrated superior performance in multimodal medical image fusion by leveraging adaptive convolution and transformer mechanisms to utilize local and global contextual information tang2022matr.
In particular, contrastive learning plays a crucial role in enhancing multimodal medical analysis under challenging conditions, such as limited data and class imbalance. For example, contrastive-based frameworks have been applied to few-shot diagnosis yin2025cognitive. In medical image classification, cong2024adaptive developed AdUniGraph, an adaptive unified contrastive learning framework with graph-based feature aggregation, which employs an adaptive loss function and convolutional graph neural networks to effectively mitigate class imbalance issues in medical imaging datasets. Furthermore, cao2025multi presented MCMCCL, a semi-supervised medical image segmentation model that integrates multi-view consistency learning and multi-scale cross-layer contrastive learning, enabling the more effective utilization of unlabeled data and enhancing segmentation accuracy. However, the greatest challenge with the approach mentioned above is that it does not consider the different characteristics of different modalities when performing modality fusion, which creates gaps in feature representation between their distributions.
To learn consistent latent representations, we designed a dual-graph architecture. The first graph utilizes patient image features extracted from a pre-trained ResNet model as node representations, where each node represents a patient. The second graph incorporates structured meta-features, as illustrated in Fig. 2 (d). We then employ a dual-structured graph attention-based fusion framework to project and scale these different modalities into a consistent feature space, as shown in Fig. 3.
@@ 2.3 Multi-Modality Fusion Strategies
3 METHODOLOGY
The schematic overview of the proposed CGMCL framework be provided in Fig. 3, illustrating how cross-modal representations are progressively aligned into a unified latent space. We propose dual-graph architecture that captures intra- and inter-modality dependencies through graph attention networks and a contrastive optimization objective to align heterogeneous features.
3.1 Problem Definition and Formulation
In clinical multimodal datasets, medical images and meta-features often have heterogeneous distributions and lack direct correspondence. To address this, we model patient similarities within each modality as separate graphs and align them through contrastive learning to improve disease classification beyond unimodal methods. Formally, given a multimodal set $\textbf{X}=\{X_{1},X_{2},...,X_{N}\}$ for $N$ patients, each sample is defined as 3-tuple $X_{i}=\{(I_{i},C_{i},Y_{i})\}_{i=1}^{N}$, where $I_{i}\in\mathbb{R}^{h\times h}$ denotes the medical image, $C_{i}\in\mathbb{R}^{F}$ the meta-features, and $Y_{i}$ corresponds to the disease labels. We first construct a non-linear model (e.g., an CNN) to generate initial feature maps $I_{i}^{\prime}$ from the images:
@@ 3.1 Problem Definition and Formulation
where the function $f(\cdot):\mathbb{R}^{D}\to\mathbb{R}^{d}$ represents the feature extractor for images, $D$ is the dimension of the initial latent space in the l1l-1 layer, and $d$ is the reduced feature dimension after the $l$-th convolution layer.
3.2 Cross-Graph Representation Construction
To overcome the limitations of conventional fusion methods, including feature concatenation and attention-based mechanisms, which tend to lose structural information among patients when heterogeneous modalities contain noise, we construct two similarity modality graphs, a structured-image graph $\mathcal{G}^{I}(\mathcal{E}^{I},\mathcal{V}^{I})$ and a structured meta-feature graph $\mathcal{G}^{C}(\mathcal{E}^{C},\mathcal{V}^{C})$, with edges $\left|\mathcal{E}^{I}\right|$ and $\left|\mathcal{E}^{C}\right|$ and vertices $\left|\mathcal{V}^{I}\right|$ and $\left|\mathcal{V}^{C}\right|$, respectively. Given the input encoded features $F_{i}$ (e.g., $I_{i}^{\prime}$ or $C_{i}$), we can construct the binary adjacency matrices $A^{I}$ and $A^{C}$ using a $K$-nearest neighbors graph pedregosa2011scikit.
$A_{ij}=\begin{cases}1&\mbox{if }F_{i}\in\mathcal{N}(F_{j})\hskip 8.0ptor\hskip 8.0ptF_{j}\in\mathcal{N}(F_{i})\\ 0&\mbox{if }F_{i}\not\in\mathcal{N}(F_{j})\hskip 8.0ptor\hskip 5.0ptF_{j}\not\in\mathcal{N}(F_{i}),\end{cases}$ (2)
@@ 3.2 Cross-Graph Representation Construction
3.3 Graph Attention Encoder
To encode both the image and clinical graphs, we adopt GAT that dynamically assigns weights to neighboring nodes by identifying the most influential patient connections during feature aggregation. Given the two input features node representations $I^{\prime}=[I^{\prime}_{1},I^{\prime}_{2},...,I^{\prime}_{N}]\in\mathbb{R}^{N\times d}$ and $C=[C_{1},C_{2},...,C_{N}]\in\mathbb{R}^{N\times F}$, we can formulate the multi-modality attention coefficients in a GAT as:
@@ 3.3 Graph Attention Encoder
$\vskip-2.84526ptH_{i}^{I^{\prime}}=\sigma\left(\sum_{j\in\mathcal{N}_{i}}\alpha^{I^{\prime}}_{ij}\mathbf{W}^{I^{\prime}}{I^{\prime}}_{j}\right),$ (7)
@@
We focus on enhancing the fusion of learned representations across cross-graph node modality embeddings. To integrate both the GAT output representations $H_{i}^{(I^{\prime})}$ and $H_{i}^{C}$ with the encoder feature information, we also incorporate a concatenation of features generated by the CNN encoder $I^{\prime}$ and the meta-features $C$ into the graph encoder. This can be expressed as:
$\mathcal{H}^{I}=\left[H_{i}^{(I^{\prime})}\mathbin{\|}I_{i}^{\prime}\right],$ (9)
@@
where $\mathbin{\|}$ denotes the concatenation operator. The matrices $\mathcal{H}^{I}\in\mathbb{R}^{N\times(F+F^{\prime})}$ and $\mathcal{H}^{C}\in\mathbb{R}^{N\times(F+F^{\prime})}$ represent the concatenated feature representations. This concatenation jointly preserves the local graph topology and the intrinsic semantic information, thereby facilitating richer multimodal feature learning.
3.4 Inter-Modality Feature Enhancement and Scaling
To address scale discrepancies between features and reduce heterogeneity modality gaps in the latent space for two single-modality output features, we introduce the inter-modality feature enhancement and scaling (IMFES) module. This module is designed to enhance the learning of intrinsic modality distributions and preserve important structural information within each modality. The process begins by applying a simple multilayer perceptron (MLP) to the concatenated tensors $\mathcal{H}^{I}$ and $\mathcal{H}^{C}$, transforming them into a single-modality probability matrix. Next, we perform element-wise multiplication between the GAT encoder outputs $H^{I^{\prime}}$ and $H^{C}$, and the non-linear MLP transforms $\zeta(\mathcal{H}^{I})$ and $\zeta(\mathcal{C}^{I})$, where $\zeta(\cdot)=\mathrm{MLP(\cdot)}$. This operation enables fine-grained extraction of individual modality features, as represented by the following equations: